# Re-ranking Using Large Language Models for Mitigating Exposure to Harmful Content on Social Media Platforms

### v1

@@ Abstract
Rajvardhan Oak1,3, Muhammad Haroon1, Claire Jo1, Magdalena Wojcieszak1, Anshuman Chhabra2 1University of California, Davis, 2University of South Florida, 3Microsoft Corporation, USA {rvoak, mharoon, wjo, mwojcieszak}@ucdavis.edu, anshumanc@usf.edu
@@ 1 Introduction
Social media platforms are powered by Machine Learning (ML) and Artificial Intelligence (AI) based recommendation algorithms and models that provide content for users. These algorithms and models are designed to maximize user engagement by learning to recommend content aligned with users’ inferred interests or traits Covington et al. (2016). However, solely optimizing for user engagement metrics can indirectly drive exposure to harmful content. For instance, a teenager interested in fitness may be recommended content promoting eating disorders, users interested in finance may encounter clickbait and scam videos, and a sad adolescent may be directed to content about depression or suicide Hilbert et al. (2024). As a result, there are serious concerns that platform recommendation systems can indirectly foment misinformation, addictions or mental health crises, and lead to other problems for individuals and society at large Haidt and Twenge (2022, 2023); Roose (2019); Tufekci (2018).
@@
In this paper, we propose methods that can circumvent both issues: that of scale as well as that of the dynamicity of harm. We utilize Large Language Models (LLMs) for this purpose as they have been shown to demonstrate stellar reasoning capabilities on natural language input in the zero-shot or few-shot learning setting Wei et al. (2022) making them ideal for harm mitigation. We find that as opposed to traditional harm classifiers such as Perspective API [^1] https://www.perspectiveapi.com/., LLMs can excel at mitigating exposure to harm in recommended content sequences simply by pairwise comparisons and re-ranking, most likely due to better utilization of their context window. The benefits of pairwise re-ranking have been explored in past work Qin et al. (2024) and we adopt this re-ranking strategy using LLMs to mitigate exposure to harmful content on platforms. This approach is visually depicted in Figure 1.
@@ 2.2 Harm Mitigation
Extensive work studies harm mitigation interventions on social media. For example, Bhargava et al. (2019) developed a tool that allows users to exert more control over their social media feeds by enabling them to consolidate and tailor content from multiple platforms. Similarly, Kovacs et al. (2018) empower users to manage their social media engagement goals via rotating time or site blockers. More germane to minimizing harmful content, other work explores ranking based interventions Celis et al. (2019); Ovadya and Thorburn (2023), such as using LLMs for curating societally beneficial recommendations. In addition, machine learning models, particularly those leveraging deep learning, are widely adopted to identify patterns indicative of harmful content, such as clickbait Ghosh et al. (2022), hate speech Mossie and Wang (2020); Del Vigna12 et al. (2017) and misinformation Wu et al. (2019); Shu et al. (2017). Recent work additionally applies LLMs to the harm detection problem  Liu et al. (2024); Ernst (2024) and also utilized machine learning toxicity detection algorithms and a browser extension that automatically hides toxic text content on users’ feed over-time Beknazar-Yuzbashev et al. (2022).
@@ 3.2 Preferential Pairwise Ranking
We present our proposed solution to the harm mitigation problem as Algorithm 1. Our proposed approach consists of a pairwise ranking component that seeks to downrank content if deemed harmful by the LLM $\mathcal{L}$. Specifically, we present $\mathcal{L}$ with pairs of content instances and query it to determine which is the relatively harmful one. Certain preference constraints ($\mathcal{C}$) determine the exact prompt used to query the LLM. To re-rank content based on relative harm, we adopt the approach from Qin et al. (2024) but modify their methodology and scoring function. For a given content instance $x$, their scoring function increments the score by 1 for every content instance deemed less relevant than $x$, and by $0.5$ for all other content instances. However, applying the scoring function as is, will also compare non-harmful content with one another, and incorrectly result in scores being higher. This can then lead to them being unfairly downranked. To address this challenge, in the pairwise ranking process, we allow the LLM $\mathcal{L}$ to decide if both instances are non-harmful, in which case there is no increment to the score. Given a content sequence $X$, we enumerate all possible pairs and compute the score for every content instance. We then re-order the content based on this score obtained.
@@ 3.3 Specifying Preference Constraints $\mathcal{C}$
Zero-Shot In the zero-shot setting, $\mathcal{C}$ asks to identify which of the two provided content instances is harmful, without explicitly specifying a definition of harm. In this approach, we utilize the LLM’s inherent understanding of harm learnt during pretraining.
Zero-Shot + Prompt Engineering We build upon the zero-shot approach by including a definition for harm in $\mathcal{C}$. We first define explicitly what we consider harmful and characteristics of harmful content, and then query $\mathcal{L}$ to identify which of the two content instances is harmful.
Few-Shot ICL We now provide representative instances of harmful content in $\mathcal{C}$, and query $\mathcal{L}$ to reason which of the two content instances is harmful based on this information. This approach is known as In-Context Learning (ICL) Dong et al. (2022).
@@ 3.5 Evaluating Re-rankings
Exponentially Weighted Normalization The Exponentially Weighted Normalization (EWN) metric provides an analytical measurement of non-harmful rankings by assigning exponentially decaying weights to ranks and then normalizing values to lie between $[0,1]$. EWN($X$) measures the goodness of the ranking in the sequence $X$ by comparing it to the best and worst ranking possible. A value of $1$ indicates that the sequence is in the optimal order and no better order can minimize exposure to harm. On the other hand, a value of 00 indicates that the sequence is in the worst possible ranking order. Following the notation described earlier, the EWN can be defined as follows:
@@ 4.1 Datasets
We employ a curated dataset of YouTube videos ($9,832$ harmful, $2,679$ harmless), which were labeled for six categories of harm: information, hate and harassment, addictive, clickbait, sexual, and physical harms. The details on the harm categories, data collection, labeling, reliability, and other aspects of this dataset are provided in  Jo et al. (2024); Jo and Wojcieszak (2025). Note that we primarily utilize this dataset for the majority of our ablation experiments because of the diversity in harm categories. We use the video descriptions as input to the LLM. To demonstrate the generalizability of our approach, we also evaluate it on two singular-harm category datasets; the Jigsaw Toxicity Classification Dataset Jigsaw (2019), which consists of comments from the Civil Comments platform, labeled for toxicity at Jigsaw, and the Measuring Hate Speech dataset D-Lab (2022) by UC Berkeley’s D-Lab, that contains annotated social media posts specifically labeled for hate speech. Additional details on each dataset are provided in Appendix B.
@@ 4.2 Content Sequences
Note that our proposed approach operates on content sequences rather than individual instances. We use the datasets described to simulate sequences of content that a user would be exposed to. We sample uniformly at random without replacement from our data to generate a sequence of $n=20$ content instances (textual data; either titles and transcripts from YouTube videos, comments, or social media posts, depending on the dataset), and generate $m=100$ such sequences to form a sequence dataset. To study the effect of the amount of harm, we generate $5$ such datasets by the varying the fraction of harmful content from $10\%$ to $50\%$ for the YouTube dataset. For the other datasets, we fix the fraction of harmful content to $30\%$ so as to reflect the typical harm ratio observed in the wild Beata Mostafavi (2020).
@@ 5.3 Varying Exemplars for Few-Shot ICL
Because the performance of ICL will depend on exemplars, we experimented with various values for the number of exemplars $N$. We modify the few-shot ICL approach and implement it in five versions by setting the value of $N$ as one of $4,8,12,16$ and $20$. We perform K-Means clustering on the harm samples (setting $K=N$) and choose a representative sample from each cluster. We conduct our analysis in the $30\%$ setting as that is the typical harm ratio observed Beata Mostafavi (2020) online. Table 7 and Figure 3 show the performance of ICL-based prompting with varying number of exemplars. Surprisingly, the performance does not necessarily improve with increasing number of exemplars; we see that EWN is the best for $N=4$ and decreases with increasing $N$. This is likely due to overfitting the LLM and the bias introduced because of the exemplars. Nevertheless, ICL still remains a better approach as compared to zero-shot or zero-shot with prompt engineering.
@@ 5.5 Experiments across LLMs
Figure 4: EWN values for Llama2, Mistral, and GPT-3.5.
@@ 6.1 Key Takeaways
Robustness to Concept Drift. Despite the differences in content types and the diversity of harm categories, our LLM approach consistently outperforms traditional methods, suggesting that the method is versatile and can be applied across diverse social media platforms and content/harm moderation challenges. Due to extensive pre-training, LLMs can generalize across various types of harm without needing explicit fine-tuning or labeled examples for each new harm type. Consequently, our approach excels at identifying and mitigating a wide range of harmful content, whether it involves misinformation, hate speech, clickbait, or addictive material, among other categories. While LLMs are still restricted temporally by training data, they can adapt to novel scenarios without requiring continuous retraining, and generalize better than traditional supervised learning methods.
@@ 7 Conclusion
In sum, we present a novel approach to mitigating exposure of harmful content on social media platforms by leveraging Large Language Models (LLMs) for re-ranking recommendation sequences. LLM-based re-ranking not only outperforms traditional classifier-based commercial moderation systems, such as the Perspective API and OpenAI Moderation API, but also exhibits robustness across various harm scenarios and content types. By utilizing zero-shot, few-shot, and prompt-engineered strategies, our method effectively generalizes to multiple forms of harm without the need for extensive labeled data, addressing both scalability and dynamicity issues inherent in content moderation. The proposed re-ranking technique shows substantial promise in reducing the likelihood of users encountering harmful content, even as the volume of such content increases in users’ social media ecosystem (as is the case for the heavy consumers of misinformation, violence, among other harmful contents). For future work, we aim to explore multi-modal input/models to further augment performance and apply our methods to other novel problem domains where content ranking plays a critical role in ensuring user safety and trust.
@@
While our proposed approach outperforms existing baselines, it has certain limitations as well. The approach involves queries to LLMs, which may be expensive in terms of time and infrastructure cost when considered at the scale that social media platforms operate (although most popular social media platforms operate with infrastructure that can support such intervention strategies [^7] https://www.facebook.com/metaai/). For smaller organizations, these issues can be minimized by lightweight (e.g. quantized) high-performance LLMs running locally (e.g. Mistral). Another limitation is modality; we only consider textual input in our algorithm and content sequences, even though visual information (e.g. video frames) could be incorporated from YouTube videos to improve performance further. We defer the study of multi-modal LLMs for re-ranking to future work.
@@
Through this work, we aim to demonstrate how LLMs can be used to minimize user exposure to harmful content online. Our work does not conduct research with human subjects directly. We do not release any new data; all data used are public, and have reported following required ethics guidelines in the collection phase. Sharing user content with third-party LLMs may introduce privacy challenges, and concerns that the data may be used for training without consent. However, social media companies can run open-source LLMs (which, from our findings, have performance comparable to proprietary ones) without relying on third parties, which would ensure that user content does not leave the company. Finally, while the primary intent of our approach is to mitigate exposure to harmful content, there exists the risk of over-moderation; LLMs, despite their advanced reasoning capabilities, might reflect inherent biases from their training data, potentially leading to the marginalization of minority perspectives or controversial opinions. Transparent moderation guidelines, diverse datasets and a human-in-the-loop moderation mechanisms can help safeguard freedom of speech and enforce effective moderation at the same time.
@@ B.1 YouTube Harms Dataset
The YouTube Harms Dataset Jo et al. (2024); Jo and Wojcieszak (2025) consists of metadata (title, transcript, descriptions) from $19,000$ YouTube videos. These were systemically collected using three approaches (i.e., keyword-based, channel-based, and external dataset integration). The videos were labeled by domain experts and crowdworkers as harmful or belonging to one or more categories of harm. The six harm categories include: information harm, hate and harassment, addictive, clickbait, sexual, and physical harms (developed based on prior work and platform guidelines). The details on the harm categories, data collection, labeling, reliability, and other aspects of this dataset are provided in  Jo et al. (2024). The data collection and annotation was IRB-approved and followed the best practices so as to minimize harm to human subjects.
@@ Appendix E Code and Reproducibility
Our code and the data used are available at the following repository: https://anonymous.4open.science/r/harm-ranking-llm/.

### v3

@@ Abstract
Rajvardhan Oak1,3, Muhammad Haroon1, Wonjeong Jo1, Magdalena Wojcieszak1, Anshuman Chhabra2 1University of California, Davis, 2University of South Florida, 3Microsoft Corporation, USA {rvoak, mharoon, wjo, mwojcieszak}@ucdavis.edu, anshumanc@usf.edu
@@ 1 Introduction
Social media platforms are powered by Machine Learning (ML) and Artificial Intelligence (AI) based recommendation algorithms and models that provide content for users. These algorithms and models are designed to maximize user engagement by learning to recommend content aligned with users’ inferred interests or traits Covington et al. (2016). However, solely optimizing for user engagement metrics can indirectly drive exposure to harmful content. For instance, a teenager interested in fitness may be recommended content promoting eating disorders, users interested in finance may encounter clickbait and scam videos, and a sad adolescent may be directed to content about depression or suicide.As a result, there are serious concerns that platform recommendation systems can indirectly foment misinformation, addictions or mental health crises, and lead to other problems for individuals and society at large Haidt and Twenge (2022, 2023); Roose (2019); Tufekci (2018).
@@
In this paper, we propose methods that can circumvent both issues: that of scale as well as that of the dynamicity of harm. We utilize Large Language Models (LLMs) for this purpose as they have been shown to demonstrate stellar reasoning capabilities on natural language input in the zero-shot or few-shot learning setting Wei et al. (2022) making them ideal for harm mitigation. We find that, as opposed to traditional harm classifiers such as Perspective API [^1] https://www.perspectiveapi.com/., LLMs can excel at mitigating exposure to harm in recommended content sequences simply by pairwise comparisons and re-ranking, most likely due to better utilization of their context window. The benefits of pairwise re-ranking have been explored in past work Qin et al. (2024) and we adopt this re-ranking strategy using LLMs to mitigate exposure to harmful content on platforms. This approach is visually depicted in Figure 1.
@@ 2.2 Harm Mitigation
Extensive work studies harm mitigation interventions on social media platforms. For example, Bhargava et al. (2019) developed a tool that allows users to exert more control over their social media feeds by enabling them to consolidate and tailor content from multiple platforms. Similarly, Kovacs et al. (2018) empower users to manage their social media engagement goals via rotating time or site blockers. More germane to minimizing harmful content, some studies explore ranking-based interventions Celis et al. (2019); Ovadya and Thorburn (2023), such as using LLMs for curating societally beneficial recommendations. In addition, machine learning models, particularly those leveraging deep learning, are often adopted to identify patterns indicative of harmful content, such as clickbait Ghosh et al. (2022), hate speech Mossie and Wang (2020); Del Vigna12 et al. (2017) and misinformation Wu et al. (2019); Shu et al. (2017). Recent work additionally applies LLMs to the harm detection problem  Liu et al. (2024); Bonagiri et al. (2025); Ernst (2024) and also utilizes machine learning toxicity detection algorithms within a browser extension to automatically hide toxic text content on users’ feed over-time Beknazar-Yuzbashev et al. (2022).
@@ 3.2 Preferential Pairwise Ranking
We present our proposed solution to the harm mitigation problem as Algorithm 1. Our proposed approach consists of a pairwise ranking component that seeks to downrank content if deemed harmful by the LLM $\mathcal{L}$. Specifically, we present $\mathcal{L}$ with pairs of content instances and query it to determine which is the relatively harmful one. Certain preference constraints ($\mathcal{C}$) determine the exact prompt used to query the LLM. To re-rank content based on relative harm, we adopt the approach from Qin et al. (2024) but modify their methodology and scoring function. For a given content instance $x$, their scoring function increments the score by 1 for every content instance deemed less relevant than $x$, and by $0.5$ for all other content instances. However, applying the scoring function as is will also compare non-harmful content with one another and incorrectly result in scores being higher. This can then lead to them being unfairly downranked. To address this challenge, in the pairwise ranking process, we allow the LLM $\mathcal{L}$ to decide if both instances are non-harmful, in which case there is no increment to the score. Given a content sequence $X$, we enumerate all possible pairs and compute the score for every content instance. We then re-order the content based on this score obtained.
@@ 3.3 Specifying Preference Constraints $\mathcal{C}$
Zero-Shot: In the zero-shot setting, $\mathcal{C}$ asks to identify which of the two provided content instances is harmful, without explicitly specifying a definition of harm. In this approach, we utilize the LLM’s inherent understanding of harm learnt during pretraining.
Zero-Shot + Prompt Engineering: We build upon the zero-shot approach by including a definition for harm in $\mathcal{C}$. We first define explicitly what we consider harmful and characteristics of harmful content, and then query $\mathcal{L}$ to identify which of the two content instances is harmful.
Few-Shot ICL: We now provide representative instances of harmful content in $\mathcal{C}$, and query $\mathcal{L}$ to reason which of the two content instances is harmful based on this information. This approach is known as In-Context Learning (ICL) Dong et al. (2022); Brown et al. (2020); Askari et al. (2025b).
@@ 3.5 Evaluating Re-rankings
Exponentially Weighted Normalization: The Exponentially Weighted Normalization (EWN) metric provides an analytical measurement of non-harmful rankings by assigning exponentially decaying weights to ranks and then normalizing values to lie between $[0,1]$. EWN($X$) measures the goodness of the ranking in the sequence $X$ by comparing it to the best and worst ranking possible. A value of $1$ indicates that the sequence is in the optimal order and no better order can minimize exposure to harm. On the other hand, a value of 00 indicates that the sequence is in the worst possible ranking order. Following the notation described earlier, the EWN can be defined as follows:
@@ 4.1 Datasets
We employ a curated dataset of YouTube videos ($9,832$ harmful, $2,679$ harmless), which were labeled for six categories of harm: information, hate and harassment, addictive, clickbait, sexual, and physical harms. The details on the harm categories, data collection, labeling, reliability, and other aspects of this dataset are provided in  Jo et al. (2024); Jo and Wojcieszak (2025). Note that we primarily utilize this dataset for the majority of our ablation experiments because of the diversity in harm categories. We use the video descriptions as input to the LLM. To demonstrate the generalizability of our approach, we also evaluate it on two singular-harm category datasets; the Jigsaw Toxicity Classification Dataset Jigsaw (2019), which contains comments from the Civil Comments platform labeled for toxicity at Jigsaw, and the Measuring Hate Speech dataset D-Lab (2022) by UC Berkeley’s D-Lab, that contains annotated social media posts specifically labeled for hate speech. Additional details on each dataset are provided in Appendix B.
@@ 4.2 Content Sequences
Note that our proposed approach operates on content sequences rather than individual instances. We use the datasets described to simulate sequences of content that a user would be exposed to. We sample uniformly at random without replacement from our data to generate a sequence of $n=20$ content instances (textual data; either titles and transcripts from YouTube videos, comments, or social media posts, depending on the dataset), and generate $m=100$ such sequences to form a sequence dataset. To study the effect of the amount of harm, we generate $5$ such datasets by the varying the fraction of harmful content from $10\%$ to $50\%$ for the YouTube dataset. For the other datasets, we fix the fraction of harmful content to $30\%$ so as to reflect the typical harm ratio observed in the wild, as determined by user surveys  Beata Mostafavi (2020).
@@ 5.3 Varying Exemplars for Few-Shot ICL
Because the performance of ICL will depend on exemplars, we experimented with various values for the number of exemplars $N$. We modify the few-shot ICL approach and implement it in five versions by setting the value of $N$ as one of $4,8,12,16$ and $20$. We perform K-Means clustering on the harm samples (setting $K=N$) and choose a representative sample from each cluster. We conduct our analysis in the $30\%$ setting  Beata Mostafavi (2020). Table 7 and Figure 3 show the performance of ICL-based prompting with varying number of exemplars. Surprisingly, the performance does not necessarily improve with increasing number of exemplars; we see that EWN is the best for $N=4$ and decreases with increasing $N$. This is likely due to overfitting the LLM and the bias introduced because of the exemplars. Nevertheless, ICL still remains a better approach as compared to zero-shot or zero-shot with prompt engineering.
@@ 5.5 Experiments across LLMs
Figure 4: EWN values for Llama2, Mistral, and GPT-3.5-Turbo.
@@ 6 Discussion
Robustness to Concept Drift. Despite the differences in content types and the diversity of harm categories, our LLM approach consistently outperforms traditional methods, suggesting that the method is versatile and can be applied across diverse social media platforms and content/harm moderation challenges. Due to extensive pre-training, LLMs can generalize across various types of harm without needing explicit fine-tuning or labeled examples for each new harm type. Consequently, our approach excels at identifying and mitigating a wide range of harmful content, whether it involves hate speech, clickbait, hate and harrassment, or addictive material, among other categories. While LLMs are still restricted temporally by training data, they can adapt to novel scenarios without requiring continuous retraining, and generalize better than traditional supervised learning methods.
@@ 7 Conclusion
In sum, we present a novel approach to mitigating exposure to harmful content on social media platforms by leveraging Large Language Models (LLMs) for re-ranking recommendation sequences. LLM-based re-ranking not only outperforms traditional classifier-based commercial moderation systems, such as the Perspective API and OpenAI Moderation API, but also exhibits robustness across various harm scenarios and content types. By utilizing zero-shot, few-shot, and prompt-engineered strategies, our method effectively generalizes to multiple forms of harm without the need for extensive labeled data, addressing both scalability and dynamicity issues inherent in content moderation. The proposed re-ranking technique shows substantial promise in reducing the likelihood of users encountering harmful content, even as the volume of such content increases in users’ social media ecosystem (as is the case for the heavy consumers of misinformation, violence, among other harmful contents). For future work, we aim to explore multi-modal input/models to further augment performance and apply our methods to other novel problem domains where content ranking plays a critical role in ensuring user safety and trust.
@@
While our proposed approach outperforms existing baselines, it has certain limitations as well. The approach involves queries to LLMs, which may be expensive in terms of time and infrastructure cost when considered at the scale that social media platforms operate (although most popular social media platforms operate with infrastructure that can support such intervention strategies [^7] https://www.facebook.com/metaai/). For smaller organizations, these issues can be minimized by lightweight (e.g. quantized) high-performance LLMs running locally (e.g. Mistral). Another limitation is modality; we only consider textual input in our algorithm and content sequences, even though visual information (e.g. video frames) could be incorporated from YouTube videos to improve performance further. We defer the study of multi-modal LLMs for re-ranking to future work. Finally, LLMs themselves can possess robustness issues Askari et al. (2025a); Chhabra et al. (2024) in downstream tasks (such as re-ranking), although these issues are exhibited by proprietary harm classification models as well Achara and Chhabra (2025).
@@
Through this work, we aim to demonstrate how LLMs can be used to minimize user exposure to harmful content online. Our work does not conduct research with human subjects directly. We do not release any new data; all data used are public, and the authors state they followed required ethics guidelines in the collection phase. Sharing user content with third-party LLMs may introduce privacy challenges, and concerns that the data may be used for training without consent. However, social media companies can run open-source LLMs (which, from our findings, have performance comparable to proprietary ones) without relying on third parties, which would ensure that user content does not leave the company. Finally, while the primary intent of our approach is to mitigate exposure to harmful content, there exists the risk of over-moderation; LLMs, despite their advanced reasoning capabilities, might reflect inherent biases from their training data, potentially leading to the marginalization of minority perspectives or controversial opinions. Transparent moderation guidelines, diverse datasets and a human-in-the-loop moderation mechanisms can help safeguard freedom of speech and enforce effective moderation at the same time.
@@ B.1 YouTube Harms Dataset
The Harmful YouTube Dataset Jo and Wojcieszak (2025) consists of metadata (title, transcript, descriptions) from $19,422$ YouTube videos. These were systemically collected using three approaches (i.e., keyword-based, channel-based, and external dataset integration). The videos were labeled by domain experts and crowdworkers as harmful or belonging to one or more categories of harm. The six harm categories include: information harm, hate and harassment, addictive, clickbait, sexual, and physical harms (developed based on prior work and platform guidelines). The details on the harm categories, data collection, labeling, reliability, and other aspects of this dataset are provided in  Jo et al. (2024); Jo and Wojcieszak (2025). The data collection and annotation was IRB-approved and followed the best practices so as to minimize harm to human subjects.
@@ Appendix E Code and Reproducibility
Our code and the data used are available at the following repository: https://github.com/rvoak/harm-ranking-llm/.